Stock in CDB Industries has a beta of 1.05. The market risk premium is 7 percent, and T-bills are currently yielding 4.0 percent. CDB’s most recent dividend was $2.90 per share, and dividends are expected to grow at a 5 percent annual rate indefinitely.

Required:

If the stock sells for $51 per share, what is your best estimate of CDB’s cost of equity? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).)

  Cost of equity %

"show steps please"
Answer #1

Using the CAPM, we find:

re = Risk-free rate + Beta(Market risk premium) = 0.04 + 1.05(0.07) = 0.1135 or 11.35%

And using the dividend growth model, the cost of equity is:

re = [D0(1 + g) / P0] + g = [$2.90(1.05)/$51] + 0.05 = 0.1097 or 10.97%

We will use the average of the two, so:

re = (0.1135 + 0.1097) / 2 = 0.1116 or 11.16%